solve the following                                2log3(x-3/x-7)+1= log3(x-3/x-1)

Grade:12th pass

1 Answers

Shreya
5 years ago
2log3(x-3/x-7) + 1= log3(x-3/x-1)
Sol.    log3(x-3/x-7)2 - log3(x-3/x-1) = -1
           log3{(x-3)(x-1)/(x-7)2}= -1
           Now we get,
           3-1={(x-3)(x-1)/(x-7)2}
          1/3= x2-4x+3/x2-14x+49
          3x2-12x+9= x2-14x+49
          2x2+2x-40=0
         x2+x-20=0
         (x+5)(x-4)=0
         x= -5 or 4(rejected as the log doesn't accept negative numbers.)
        Ans is x=-5